OpenStudy (anonymous):

Tracy needs the to find the length and width of the equipment storage room where she works. The room is 3 meters longer than it is wide. Its perimeter is 18 meters. Find the width in meters.

OpenStudy (anonymous):

Assume width= x Therefore since length is 3m more than width,x, what do u think would be the length? (hint: its an expression ie, no equal sign)

OpenStudy (anonymous):

2? @kalahkid

OpenStudy (anonymous):

No, not quite.. it would be x + 3 Remember since we dont know the exact igure we're letting the width be x

OpenStudy (anonymous):

And since the length is 3m longer it means we add 3 to the width x. So, we get x+3

OpenStudy (anonymous):

Understand so far @atlantaslast ?

OpenStudy (anonymous):

Yeah I understand

OpenStudy (anonymous):

okay so what's the formula for perimeter of a rectangle?

OpenStudy (anonymous):

p=2(l+w)

OpenStudy (anonymous):

2l+2w basically

OpenStudy (anonymous):

Good so just substitute what we got for length and width

OpenStudy (anonymous):

So what would it be? Keep in mind they gave you the perimeter 18m

OpenStudy (anonymous):

2l+3w=18?

OpenStudy (anonymous):

Good but remember l= x+3 and w = x, so put that in

OpenStudy (anonymous):

im a bit confused

OpenStudy (anonymous):

the length,l, in this case is x+3 and the width,w, is x. They gave us this information plus perimeter 18m and we have to use this to find the width. What's confusing though?

OpenStudy (anonymous):

Perimeter,P=18 l = x+3 w=x Formula to be used : 2l + 2w ---> 2(x+3) + 2(x)=18

OpenStudy (anonymous):

I was going to put that equation down but I wasn't entirely sure but now I get it so it would be 2x+6+2x=18, so if we add the like terms it would be 4x+6=18 right?

OpenStudy (anonymous):

@kalahkid

OpenStudy (kevin8767):

Let the width of the room = x m then the length = (x+3) m Perimeter of the room = 2(Length+Width) 2(x+3+x) = 18 4x+6 = 18 4x = 18-6 = 12 x = 12/4 = 3 Therefore the width of the room = 3 meters that's the correct width of the room